What is the volume of the largest rectangular parallelepiped which can be inscribed in the ellipsoid x2/9+y2/16+z2/36=1? Show that the answer you get gives you the largest volume.(DO NOT USE LAGRANGE MULTIPLIERS)


1
Expert's answer
2021-12-05T17:41:54-0500

Let P=(x,y,z) be a point on the ellipsoid with x,y,z>0. Take the eight different points with

Pi​(±xyz)

These points are the vertices of a parallelepiped with the side length

2x,2y and 2z

Then, the volume parallelepiped is:

"V=2x\u22c52y\u22c52z=8xyz"

"z=6\sqrt{1-x^2/9-y^2/16}"


"V=48xy\sqrt{1-x^2/9-y^2/16}"


"V_x=48y\sqrt{1-x^2/9-y^2/16}-\frac{16x^2y}{3\sqrt{1-x^2/9-y^2/16}}=0"


"9(1-x^2\/9-y^2\/16)-x^2=0"


"V_y=48x\sqrt{1-x^2/9-y^2/16}-\frac{16y^2x}{3\sqrt{1-x^2/9-y^2/16}}=0"


"9(1-x^2\/9-y^2\/16)-y^2=0"


"x^2=y^2"

"x=y"

"9-2x^2-9x^2\/16=0"

"41x^2=144"

"x=y=\sqrt{41}/12"


"V_{xx}=48y\sqrt{1-x^2/9-y^2/16}-\frac{16x^2y}{3\sqrt{1-x^2/9-y^2/16}}="


"=-\frac{32xy}{3\sqrt{1-x^2/9-y^2/16}}"


"z=6\sqrt{1-x^2/9-x^2/16}=\sqrt{144-25x^2}/2=\sqrt{19711}/24"


"V_{max}=\frac{8\cdot41\sqrt{19711}}{144\cdot24}=13.32"